A standard six sided die is rolled 11 times. What is the expected value of the number of times an odd number will be rolled. Do not round your answer.
Answer #1

Probability of odd number rolled = 3 / 6 = 1/2

n = 11

P = 1/2

We can find the expected value by Binomial distribution

X B ( n = 11 , P = 0.5 )


(Expected value ) Mean = n * P = ( 11 * 0.5 ) = 5.5